# Differentially Private Clustered Federated Learning

### v5

@@ 3 Definitions, Notations and assumptions
        // full batch sizes in the first round
@@ 4       $z_{i}\leftarrow$RDP($\epsilon,\delta,b_{i}^{1},b_{i}^{>1},N_{i},K,E$)
                   $\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}\leftarrow$DPSGD ($\mathbf{\bm{\theta}}^{init},b_{i}^{1},N_{i},K,z_{i},c$) ;
                    // DPSGD with full batch size on $\bm{\theta}^{init}$
10
11
@@
                   $M=\operatorname*{arg\,max}_{M^{\prime}}\texttt{MSS}\Big{(}\textbf{{GMM}}(\Delta\tilde{\mathbf{\bm{\theta}}}_{1}^{1},\ldots,\Delta\tilde{\mathbf{\bm{\theta}}}_{n}^{1};M^{\prime})\Big{)}$ ;
                    // when $M$ is unknown (Section 4.4)
14
@@
            $\{\pi_{1},\ldots,\pi_{n},\texttt{MPO}\}=\textbf{{GMM}}(\Delta\tilde{\mathbf{\bm{\theta}}}_{1}^{1},\ldots,\Delta\tilde{\mathbf{\bm{\theta}}}_{n}^{1};M)$ ;
              // 1st stage: GMM with $M$ components
@@
              // $E_{c}$ is set based on MPO (Section 4.4)
17
            Initialize: $\bm{\theta}_{1}^{2}=\ldots=\bm{\theta}_{M}^{2}=\mathbf{\bm{\theta}}^{\textit{init}}$ ;
              // initialize the $M$ cluster models uniformly
@@
                    // 3rd stage: "local" clustering
29
@@
              // DPSGD with batch size $b_{i}^{>1}$ on $\mathbf{\bm{\theta}}_{R^{e}(i)}^{e}$
34
@@ 38             $w_{i}^{e}\leftarrow\frac{1}{\sum_{j=1}^{n}\mathds{1}_{R^{e}(j)=R^{e}(i)}}$
41
            $\mathbf{\bm{\theta}}_{m}^{e+1}\leftarrow\mathbf{\bm{\theta}}_{m}^{e}+\sum_{i\in\{1,\dots,n\}}\mathds{1}_{R^{e}(i)=m}w_{i}^{e}\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{e}$ ;
              // $i$ contributes to $R^{e}(i)$
@@
43
@@ 4 Methodology and proposed algorithm
Figure 2: PCA visualization of updates $\{\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}\}_{i=1}^{n}$ on 2D space. Left: $\epsilon_{i}=10$, $b_{i}^{e}=32$ for all $i$ and $e$. Right: $\epsilon_{i}=10$, $b_{i}^{1}=b^{1}=N=6600$, i.e., full batch sizes (assuming $N_{i}=N=6600$ for all clients), and $b_{i}^{>1}=32$ for all $i$. $\mu_{m}^{*}(b^{1})$ shows the center of cluster $m$, when all clients use batch size $b^{1}$ during the first round. The model updates are obtained from clients running DPSGD for $K=1$ epochs locally on CIFAR10 with covariate shift (rotation) across clusters, and under the same values as in Figure 3.
@@ 4.1 R-DPCFL algorithm
In Sections 4.2 and 4.3, we provide theoretical justifications and analysis for the proposed method, especially the usage of full batch size in the first round.
@@
The DP noise in $\{\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}\}_{i=1}^{n}$ makes it harder for the server to cluster clients by learning a GMM on the model updates. The following lemma quantifies this noise amount by extending a result in (Malekmohammadi et al., 2024) to when different batch sizes are used in the first and the next rounds:
@@
Let us assume $\bm{\theta}_{i}^{e,0}$ is the model parameter passed to client $i$ at the beginning of round $e$. After $K$ local epochs with step size $\eta_{l}$, the client generates the noisy DP model update $\Delta\tilde{\bm{\theta}}_{i}^{e}(b_{i}^{e})$ at the end of the round. The amount of noise in the resulting model update can be found as:
@@
The first conclusion from the lemma is that the noise level in $\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{e}$ rapidly declines as $b_{i}^{e}$ increases: See Figure 3 and the effect of batch size $b_{i}^{1}$ on $\texttt{Var}(\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}|\mathbf{\bm{\theta}}_{i}^{init})$ (on the left) and the effect of batch size $b_{i}^{>1}$ on $\texttt{Var}(\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{e}|\mathbf{\bm{\theta}}_{i}^{e,0})$ ($e>1$) (on the right). Let us consider $e=1$ especially: If all clients use full batch sizes in the first round (i.e., $b_{i}^{1}=N_{i}$ for every client $i$), it becomes much easier for the server to cluster them at the end of the first round by learning a GMM on $\{\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}\}_{i=1}^{n}$, as their updates become more separable. An illustration of this is shown in Figure 2. In the next section, we will provide a theoretical justification for this observation. As the second key takeaway, Figure 3, left shows that in order to make $\{\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}\}_{i=1}^{n}$ less noisy, we have to make $\{b_{i}^{1}\}_{i=1}^{n}$ as large as possible and also keep $\{b_{i}^{>1}\}_{i=1}^{n}$ small [^2] In fact, there is a close relation between the result of 4.1 and the law of large numbers. See Appendix H for more details.
@@ 4.2.1 Effect of batch sizes $\{b_{i}^{1}\}_{i=1}^{n}$ on the separation between clusters
In this section, we provide a theoretical justification for the observation in Figure 2, right. For simplicity, let us assume clients have the same dataset sizes and first batch sizes: $\forall i:N_{i}=N,b_{i}^{1}=b^{1}$. Also, remember that $\mathbf{\bm{\theta}}_{i}^{1,0}=\mathbf{\bm{\theta}}^{\textit{init}}$. Having uniform privacy parameters $(\epsilon,\delta)$, we have: $\forall i:~{}\sigma_{i}^{1^{2}}(b^{1}):=\texttt{Var}[\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}(b^{1})|\mathbf{\bm{\theta}}^{\textit{init}}]=\sigma^{1^{2}}(b^{1})$. Hence, we can consider the model updates $\{\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}(b^{1})\}_{i=1}^{n}$ as the samples from a mixture of $M$ Gaussians with mean, covariance matrix, prior probability parameters: $\psi^{*}(b^{1})=\{\mu_{m}^{*}(b^{1}),\Sigma_{m}^{*}(b^{1}),\alpha_{m}^{*}\}_{m=1}^{M}$, where $\forall m:\alpha_{m}^{*}>0$ and $\mu_{m}^{*}(b^{1})\neq\mu_{m^{\prime}}^{*}(b^{1})$ ($m\neq m^{\prime}$), due to data heterogeneity:
@@ 5 Evaluation
Datasets, models and baseline algorithms: We evaluate our proposed method on three benchamark datasets, including: MNIST (Deng, 2012), FMNIST (Xiao et al., 2017) and CIFAR10 (Krizhevsky, 2009), with heterogeneous data distributions from covariate shift (rotation; $P_{i}(x)$ varies across clusters) (Kairouz et al., 2021; Werner et al., 2023) and concept shift (label flip; $P_{i}(y|x)$ varies across clusters) (Werner et al., 2023), which are the commonly used data splits for clustered FL (see Appendix D). We consider four clusters of clients indexed by $m\in\{0,1,2,3\}$ with $\{3,6,6,6\}$ clients, where the smallest cluster is considered as the minority cluster. We compare our method with most recent related DPFL algorithms under an equal total sample-level privacy budget $\epsilon$: 1. Global (Noble et al., 2021): clients run DPSGD locally and send their model updates to the server for aggregation and learning one global model 2. Local (Liu et al., 2022a): clients do not participate FL and learn a local model by running DPSGD on their local data 3. A DP extension of IFCA (Ghosh et al., 2020; Liu et al., 2022a): local loss/accuracy-based clustering performed by clients on existing cluster models 4. MR-MTL (Liu et al., 2022a): uses model personalization to learn one model for each client 5. O-DPCFL: an oracle algorithm which has the knowledge of the true clusters from the first round. For R-DPCFL and IFCA, we use exponential mechanism (Rogers & Steinke, 2021), which satisfies zero concentrated DP (z-CDP) (Bun & Steinke, 2016), to privatize clients’ local cluster selections.
@@ 5.1 Results
RQ2: How does the minority cluster benefit from R-DPCFL? Figure 5 compares different algorithms based on the average test accuracy of the clients belonging to the minority cluster. R-DPCFL leads to a better overall performance for the minority clients, by virtue of its correct and robust cluster detection. Correct detection of the minority cluster prevents it from getting mixed with other majority clusters and leads to a utility improvement for its clients. In contrast, IFCA has a lower success rate in detecting the minority cluster (Figure 5, bottom row) and provides a lower overall performance for them. Similarly, Local and MR-MTL lead to a low performance for the minority, as they are conditioned on a global model that is learned from federation across all clients and provides a low performance for the minority. Detecting and improving the performance of minority clusters is important, as the different clusters in DP clustered FL systems may not be of the same size, and failure in detecting them correctly leads to low performance for the ones with smaller sizes.
@@ 6 Conclusion
We proposed a DP clustered FL algorithm, which addresses sample-level privacy in FL systems with structured data heterogeneity. By clustering clients based on both their model updates and training loss/accuracy values, and mitigating noise impacts with large initial batch sizes, our approach enhances clustering accuracy and mitigates DP’s disparate impact on utility, all with minimal computational overhead. Moreover, the robustness to noise and easy parameter selection of the proposed approach shows its applicability to DP clustered FL settings. While envisioned for DPFL systems with large local datasets, the method is capable of compensating for moderate dataset sizes by using smaller batch sizes after the first round. In the future, we aim to extend this approach to be suitable for scarce data scenarios, which might be found in cross-device FL settings.
7 Acknowledgement
Funding support for project activities has been partially provided by Canada CIFAR AI Chair, Facebook award, Google award, and MEI award. We also express our gratitude to Compute Canada for their support in providing facilities for our evaluations.
@@ C.1 Renyi Differential Privacy (RDP)
A randomized mechanism $\mathcal{M}:\mathcal{A}\to\mathcal{R}$ with domain $\mathcal{D}$ and range $\mathcal{R}$ satisfies $(\alpha,\epsilon)$ RDP with order $\alpha>1$ if for any two adjacent inputs $\mathcal{D}$, $\mathcal{D}^{\prime}\in\mathcal{A}$, which differ only by a single record,
@@
For $\alpha=1$, we have $D_{1}(P||Q):=\mathbb{E}_{x\sim p}\bigg{[}\log\big{(}\frac{P(x)}{Q(x)}\big{)}\bigg{]}$, which is the KL divergence between $P$ and $Q$. RDP can be used for composition of private mechanisms: if an algorithm has $E$ steps and each step satisfies $(\alpha,\epsilon)$-RDP, the algorithm satisfies $(\alpha,E\epsilon)$-RDP. RDP can also be used for composition of heterogeneous private mechanisms, e.g., for accounting privacy of R-DPCFL, which uses different batch sizes in the first and next rounds. The following lemma is about conversion of $(\alpha,\epsilon)$-RDP to standard $(\epsilon,\delta)$-DP (Definition 3.1).

### v6

@@ 3 Definitions, Notations and assumptions
        // full batch size
4
@@ 5      $z_{i}\leftarrow$ RDP($\epsilon,\delta,b_{i}^{1},b_{i}^{>1},N_{i},K,E$)
11                   $\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}\leftarrow$ DPSGD ($\mathbf{\bm{\theta}}^{init},b_{i}^{1},N_{i},K,z_{i},c$)
12
@@
                   $M=\operatorname*{arg\,max}_{M^{\prime}}\texttt{MSS}\Big{(}\texttt{GMM}(\Delta\tilde{\mathbf{\bm{\theta}}}_{1}^{1},\ldots,\Delta\tilde{\mathbf{\bm{\theta}}}_{n}^{1};M^{\prime})\Big{)}$ ;
                    // set $M$ (Section 4.4)
@@
            $\{\pi_{1},\ldots,\pi_{n},\texttt{MPO}\}=\texttt{GMM}(\Delta\tilde{\mathbf{\bm{\theta}}}_{1}^{1},\ldots,\Delta\tilde{\mathbf{\bm{\theta}}}_{n}^{1};M)$ ;
              // 1st stage: GMM
17
@@
              // set $E_{c}$ (Section 4.4)
@@
19            Initialize cluster models uniformly: $\bm{\theta}_{1}^{2}=\ldots=\bm{\theta}_{M}^{2}=\mathbf{\bm{\theta}}^{\textit{init}}$
@@
                    // 3rd stage: private local clustering
@@
              // batch size $b_{i}^{>1}$
@@ 39             $w_{i}^{e}\leftarrow\frac{1}{\sum_{j=1}^{n}\mathds{1}_{R^{e}(j)=R^{e}(i)}}$
43            $\mathbf{\bm{\theta}}_{m}^{e+1}\leftarrow\mathbf{\bm{\theta}}_{m}^{e}+\sum_{i\in\{1,\dots,n\}}\mathds{1}_{R^{e}(i)=m}w_{i}^{e}\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{e}$
@@ 4 Methodology and proposed algorithm
Figure 2: PCA visualization of updates $\{\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}\}_{i=1}^{n}$ on 2D space. Left: $\epsilon_{i}=10$, $b_{i}^{e}=32$ for all $i$ and $e$. Right: $\epsilon_{i}=10$, $b_{i}^{1}=b^{1}=N=6600$, i.e., full batch sizes (assuming $N_{i}=N=6600$ for all clients), and $b_{i}^{>1}=32$ for all $i$. The empty markers show the centers of the Gaussian components. The model updates are obtained from clients running DPSGD for $K=1$ epochs locally on CIFAR10 with covariate shift (rotation) across clusters, and under the same values as in Figure 3.
@@ 4.1 R-DPCFL algorithm
In Sections 4.2 and 4.3, we provide theoretical justifications for the usage of full batch size in the first round.
@@
It is the DP noise in $\{\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}\}_{i=1}^{n}$ that makes it hard for the server to cluster clients by learning a GMM on the model updates. 4.1 extends a result in (Malekmohammadi et al., 2024) and shows that the amount of noise in $\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{e}$ at the end of each round $e$ rapidly drops when $b_{i}^{e}$ increases.
@@
Let $\approx$ denote approximation. After $K$ local epochs with step size $\eta_{l}$, client $i$ generates the noisy DP model update $\Delta\tilde{\bm{\theta}}_{i}^{e}(b_{i}^{e})$ at the end of the round $e$. The amount of noise in the resulting model update can be found as:
@@
The first conclusion from the lemma is that the noise level in $\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{e}$ rapidly declines as $b_{i}^{e}$ increases: See Figure 3 and the effect of batch size $b_{i}^{1}$ on $\texttt{Var}(\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}|\mathbf{\bm{\theta}}_{i}^{init})$ (on the left); and the effect of batch size $b_{i}^{>1}$ on $\texttt{Var}(\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{e}|\mathbf{\bm{\theta}}_{i}^{e,0})$ ($e>1$) (on the right). Let us consider $e=1$ especially: If all clients use full batch sizes in the first round (i.e., $b_{i}^{1}=N_{i}$ for every client $i$), it becomes much easier for the server to cluster them at the end of the first round by learning a GMM on $\{\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}\}_{i=1}^{n}$, as their updates become more separable. An illustration of this is shown in Figure 2. As the second key takeaway, Figure 3, left shows that in order to make $\{\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}\}_{i=1}^{n}$ less noisy, we have to make $\{b_{i}^{1}\}_{i=1}^{n}$ as large as possible and also keep $\{b_{i}^{>1}\}_{i=1}^{n}$ small [^2] In fact, there is a close relation between the result of 4.1 and the law of large numbers. See Appendix H for more details. In the next section, we will provide a theoretical justification for the observation in Figure 2.
@@ 4.2.1 Effect of batch sizes $\{b_{i}^{1}\}_{i=1}^{n}$ on the separation between clusters
In order to theoretically understand the reason behind the observation in Figure 2, let us assume clients have the same dataset sizes and first batch sizes for simplicity: $\forall i:N_{i}=N,b_{i}^{1}=b^{1}$. Also, remember that $\mathbf{\bm{\theta}}_{i}^{1,0}=\mathbf{\bm{\theta}}^{\textit{init}}$. Having uniform privacy parameters $(\epsilon,\delta)$, we have: $\forall i:~{}\sigma_{i}^{1^{2}}(b^{1}):=\texttt{Var}[\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}(b^{1})|\mathbf{\bm{\theta}}^{\textit{init}}]=\sigma^{1^{2}}(b^{1})$. Hence, we can consider the model updates $\{\Delta\tilde{\mathbf{\bm{\theta}}}_{i}^{1}(b^{1})\}_{i=1}^{n}$ as the samples from a mixture of $M$ Gaussians with mean, covariance matrix, prior probability parameters: $\psi^{*}(b^{1})=\{\mu_{m}^{*}(b^{1}),\Sigma_{m}^{*}(b^{1}),\alpha_{m}^{*}\}_{m=1}^{M}$, where $\forall m:\alpha_{m}^{*}>0$ and $\mu_{m}^{*}(b^{1})\neq\mu_{m^{\prime}}^{*}(b^{1})$ ($m\neq m^{\prime}$), due to data heterogeneity:
@@ 5 Evaluation
Datasets, models and baseline algorithms: We evaluate our proposed method on three benchamark datasets, including: MNIST (Deng, 2012), FMNIST (Xiao et al., 2017) and CIFAR10 (Krizhevsky, 2009), with heterogeneous data distributions from covariate shift (rotation; $P_{i}(x)$ varies across clusters) (Kairouz et al., 2021; Werner et al., 2023) and concept shift (label flip; $P_{i}(y|x)$ varies across clusters) (Werner et al., 2023), which are the commonly used data splits for clustered FL (see Appendix D). We consider four clusters of clients indexed by $m\in\{0,1,2,3\}$ with $\{3,6,6,6\}$ clients, where the smallest cluster is considered as the minority cluster. We compare our method with most recent related DPFL algorithms under an equal total sample-level privacy budget $\epsilon$: 1. Global (Noble et al., 2021): clients run DPSGD locally and send their model updates to the server for aggregation and learning one global model 2. Local (Liu et al., 2022a): clients do not participate FL and learn a local model by running DPSGD on their local data 3. A DP extension of IFCA (Ghosh et al., 2020; Liu et al., 2022a): local loss/accuracy-based clustering performed by clients on existing cluster models 4. MR-MTL (Liu et al., 2022a): uses model personalization to learn one model for each client 5. O-DPCFL: an oracle algorithm which has the knowledge of the true clusters from the first round. For R-DPCFL and IFCA, we use exponential mechanism (Rogers & Steinke, 2021), which satisfies zero concentrated DP (z-CDP) (Bun & Steinke, 2016), to privatize clients’ local cluster selections. See also Appendix F for further experimental results.
@@ 5.1 Results
RQ2: How does the minority cluster benefit from R-DPCFL? Figure 5 compares different algorithms based on the average test accuracy of the clients belonging to the minority cluster. R-DPCFL leads to a better overall performance for the minority clients, by virtue of its correct and robust cluster detection. Correct detection of the minority cluster prevents it from getting mixed with other majority clusters and leads to a utility improvement for its clients. In contrast, IFCA has a lower success rate in detecting the minority cluster (Figure 5, bottom row) and provides a lower overall performance for them. Similarly, Local and MR-MTL lead to a low performance for the minority, as they are conditioned on a global model that is learned from federation across all clients and provides a low performance for the minority. In DP clustered FL, detecting and improving the performance of minority clusters is important, as failure in detecting them correctly leads to low performance for the clusters with smaller sizes.
@@ 6 Conclusion
We proposed a DP clustered FL algorithm, which addresses sample-level privacy in FL systems with structured data heterogeneity. By clustering clients based on both their model updates and training loss/accuracy values, and mitigating noise impacts with large initial batch sizes, our approach enhances clustering accuracy and mitigates DP’s disparate impact on utility, all with minimal computational overhead. Moreover, the robustness to noise and easy parameter selection of the proposed approach shows its applicability to DP clustered FL settings. While envisioned for DPFL systems with large local datasets, the method is capable of compensating for moderate dataset sizes by using smaller batch sizes after the first round. In the future, we aim to extend this approach to be suitable for scarce data scenarios.
@@ C.1 Renyi Differential Privacy (RDP)
A randomized mechanism $\mathcal{M}:\mathcal{A}\to\mathcal{R}$ with domain $\mathcal{D}$ and range $\mathcal{R}$ satisfies $(\alpha,\epsilon)$-RDP with order $\alpha>1$ if for any two adjacent inputs $\mathcal{D}$, $\mathcal{D}^{\prime}\in\mathcal{A}$, which differ only by a single record,
@@
For $\alpha=1$, we have $D_{1}(P||Q):=\mathbb{E}_{x\sim p}\bigg{[}\log\big{(}\frac{P(x)}{Q(x)}\big{)}\bigg{]}$, which is the KL divergence between $P$ and $Q$. RDP can be used for composition of private mechanisms (Mironov, 2017):
Proposition C.2.
Let $\mathcal{M}_{1}:\mathcal{A}\to\mathcal{R}_{1}$ be $(\alpha,\epsilon_{1})$-RDP and $\mathcal{M}_{2}:\mathcal{R}_{1}\times\mathcal{A}\to\mathcal{R}_{2}$ be $(\alpha,\epsilon_{2})$-RDP. Then, the mechanism $\mathcal{M}_{3}=(X,Y)$ defined as $X\leftarrow\mathcal{M}_{1}(\mathcal{A})$ and $Y\leftarrow\mathcal{M}_{2}(\mathcal{A},X)$ satisfies $(\alpha,\epsilon_{1}+\epsilon_{2})$-RDP.
Therefore, if an algorithm has $E$ steps and each step satisfies $(\alpha,\epsilon)$-RDP, the algorithm satisfies $(\alpha,E\epsilon)$-RDP. RDP can also be used for composition of heterogeneous private mechanisms, e.g., for accounting privacy of R-DPCFL, which uses different batch sizes in the first and the subsequent rounds. The following lemma is about conversion of $(\alpha,\epsilon)$-RDP to standard $(\epsilon,\delta)$-DP (Definition 3.1).
@@
The Gaussian mechanism satisfies $(\alpha,\epsilon)$-RDP, based on the following Proposition from (Mironov, 2017):
Proposition C.4.
If $f:\mathcal{A}\to\mathcal{R}$ has sensitivity $c$, then its randomization with a Gaussian mechanism with noise variance $\sigma_{DP}^{2}$ satisfies $(\alpha,\frac{\alpha c^{2}}{2\sigma_{DP}^{2}})$-RDP.